two angles of a triangle measure 135 degrees and 11 degrees what is the measure of the third angle?
Answer by nerdybill(5404)

two angles of a triangle measure 135 degrees and 11 degrees what is the measure of the third angle?
.
Since the sum of the three interior angles of any triangle is 180 degrees we have:
Let t = third angle
then
t + 135 + 11 = 180
t + 146 = 180
t = 180 - 146
t = 34 degrees (measure of third angle)